OpenStudy (joshiscool):

Which value of x makes the equation true? x – 18.7 = –12 x =

OpenStudy (joshiscool):

brb

OpenStudy (baitcaster13):

sry dude i suck at math

alones (alones):

x−18.7=−12 add 18.7 to both sides. x−18.7+18.7=−12+18.7 x=6.7

OpenStudy (joshiscool):

@AloneS Thanks.

alones (alones):

My pleasure ;)

OpenStudy (joshiscool):

I have one more